The chlorine-containing narcotic drug pentaerythritol chloral has 21.51% C, 2.22% H, and 17.64% O, by mass, as its other elements. Its molecular mass is 726 u. What are the empirical and molecular formulas of pentaerythritol chloral?



1
Expert's answer
2021-11-11T14:21:01-0500

% of C = 21.51

%/ atomic mass = 21.51 / 12 = 1.79

% of H = 2.22%

% / atomic mass/ 2.22 / 1 = 2.22

% of O = 17.64 %

% / atomic mass = 17.64 / 16 = 1.1

% of Cl = 100 - ( 21.51 + 2.22 + 17.64 ) = 58.63

% / atomic mass = 58.63 / 35.5 = 1.65

Dividing by least ratio to each elements ratio

For C 1.79 / 1.1 = 1.627

For H 2.22 / 1.1 = 2

For O 1.1 / 1.1 = 1

For Cl 1.65 / 1.1 = 1.5

Simplest ratio between elements

For C 1.627 × 8 = 13

For H 2 × 8= 16

For O 1× 8= 8

For Cl 1.5 × 8 = 12

Empirical formula = C13H16​​​​​O8Cl12

Empirical formula mass

= 13 ×12 + 16×1 +8×16 + 12 ×35.5 =726

Valence (n) = Molecular formula wt/ Empirical formula wt

= 726 / 726 = 1

Molecular formula = Empirical formula =

C13H16​​​​​O8Cl12